# Personality Structured Interview for Large Language Model Simulation in Personality Research

### v1

@@ 4.1 General Settings
Psychological Scale
We used Big Five Inventory-2 (BFI-2; Soto and John, 2017): The BFI-2 is designed to capture three core facets of each of the FFM personality factors. Each facet is measured by two positively worded items and two negatively worded items, resulting in 60 items in total. Human respondents and LLMs were instructed to indicate the degree to which they agree with each item on a 5-point scale (1 = “Strongly disagree”, 2 = “Somewhat disagree”, 3 = “Neither agree nor disagree”, 4 = “Somewhat agree”, 5 = “Strongly agree”). Specific scale items and scoring criteria can be found in Table 5, Table 6 and Table 7 in Appendix E .
@@
We tested two LLMs, encompassing both closed-source and open-source models: GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024), and Llama 3 (llama-3-70b-instruct) (AI@Meta, 2024). We include these models to allow for cross-validation and comparative analysis, ensuring that the findings are robust and not specific to a single model.
GPT-4o and Llama 3 are among the current leading models, making them ideal for assessing the latest advancements in LLM performance. To ensure reproducibility, the temperature of the LLMs was set to zero to generate deterministic responses.
@@ 4.2 Response Similarity
The main purpose of this experiment is to assess the similarity between responses generated by LLMs based on the PSI method and human self-reports.
@@
Because the data generated by LLMs correspond on a one-to-one basis with data from human respondents, we used the Mean Absolute Error (MAE) and Pearson correlation coefficient ($r$) to examine the strength of the similarity between them. Smaller MAE and higher $r$ indicate greater similarity. For the $r$ calculation formula, see Appendix G.  Park et al. (2024) employed a Life Interview simulation, which is currently the most information-rich method for personality simulation. It is also one of the few papers that provides a direct one-to-one performance comparison with human samples. The reported MAE and $r$ with corresponding human samples are likely the best values available so far. We used this as a standard for comparison.
If data from human respondents are regarded as the gold-standard criterion, this test can also be considered an examination of criterion-related validity, which assesses the correlation between a measure and an external standard (Cronbach and Meehl, 1955).
@@ 4.3 Human Personality Distribution Simulation
The main purpose of this experiment is to investigate further the differences among various methods of simulating human personality distributions. We compared the PSI method, the Persona method—which relies on dialogue information (Zhang et al., 2018)—and the Shape method, which uses adjective-based dimensional categorization (Serapio-García et al., 2023), to observe differences when simulating a normal human sample. The specific descriptions of the Persona method and Shape method are shown in Appendix G.
@@
The human samples used for this experiment were collected as part of a broader project related to personality assessment through Prolific and received Institutional Review Board approval. Respondents were instructed to complete a set of demographic questions, the BFI-2, and a set of criterion measures. The respondents were compensated with \$3.75 for their participation. In total, 1,559 respondents provided valid responses. On average, participants were in their early 40s (M = 42.29 years, SD = 11.79), 50.80% identifying as men, 49.20% identifying as women.
@@
To quantify the degree of similarity between human and LLM responses, we leveraged multiple metrics. At the domain and facet levels, we compared the mean (M; sample mean of each domain and facet), standard deviation (SD; sample standard deviation of each domain and facet), Cronbach’s alpha, and correlations among scores. At the item level, we also compared the mean and standard deviation of item scores (sample mean and standard deviation of each item).
Specifically, we also used MAE and $r$ to quantify similarities in these metrics. Here, however, MAE and $r$ are based on sample-level analysis rather than the one-to-one correspondence seen in the previous experiment. Our focus is on whether, when simulating an overall human sample, similarities can be observed across various measurement levels—including items, facets, and domains.
@@
Additionally, separately for human and LLM responses, we fitted a three-factor confirmatory factor analysis (CFA; Jöreskog, 1969) model (TFM; where facets are “factors”) to the responses to each domain of the BFI-2. Aside from the facet structure of each domain, we also used the facet scores as indicators and fitted the FFM. Model fit, standardized factor loadings, and latent correlations among the facets were also compared between human and LLM responses. Tucker’s congruence coefficient (TCC; Tucker, 1951) and MAE of factor loadings were used to quantify the similarity between factor solutions from human and LLM responses. For the CFA model, model fit information, and the TCC calculation formula, see Appendix G. This evaluation framework is adaptable and can be used to analyze the fidelity of other simulated psychometric data (see Appendix H for further details).
@@ 4.4 Personality-Related Behavioral Performance
The primary aim of this experiment is to explore whether LLMs, when assigned specific personality settings, exhibit behaviors theoretically aligned with those personalities. This is an exploratory study, as LLMs generate responses based on statistical probabilities derived from the training corpus (Yang et al., 2024), making it uncertain whether assigning personality settings (through the PSI data) will influence the model to act consistently with that personality. Fortunately, the data we have collected through the PSI method includes LLM ratings of human behaviors, providing a valuable basis to test this hypothesis.
@@
We collected self-report data on two classic types of workplace behaviors: organizational citizenship behavior (OCB) and counterproductive work behavior (CWB). Both types of behaviors are widely supported by research as being related to personality (e.g., Organ and Ryan, 1995; Berry et al., 2007).
These data were collected using the scale developed by Spector and Fox (2010) and prompted the LLM to respond to the same questions (for a detailed description of Spector and Fox’s (2010) measures, see Appendix G; for specific prompt details, refer to Appendix F).
@@
We primarily focus on the $r$ between personality domains from different data sources and OCB, as well as CWB. We anticipate that the $r$ between self-reported personality domains and OCB/CWB in LLM simulation will closely align with those observed in the human participants.
5 Experimental Results
Here, we present the results of the three experiments mentioned above: Response Similarity (§5.1), Human Personality Distribution Simulation (§5.2), and Personality-Related Behavioral Performance (§ 5.3).
5.1 Response Similarity Results
Table 1 provides a clear illustration of the similarity between the personality generated using the PSI method and the corresponding human self-reported personality data. The results show that regardless of the model used, the average correlation is around .5, suggesting a moderately strong positive relationship and a significant association between the two variables. In particular, compared to the Life Interview method (Park et al., 2024), the PSI method consistently outperforms or matches the Life Interview approach across nearly all metrics, as indicated by its similar or lower MAE and similar or higher $r$.
Notably, the Life Interview method relies on up to two hours of interview data, while the PSI method achieves comparable results using interview data consisting of just 32 questions, with an average duration of about 34 minutes. This further underscores the advantage of the PSI method in effectively leveraging LLMs to simulate personality profiles.
Domain MAE $r$
PSI GPT-4o PSI Llama3 Life Interview PSI GPT-4o PSI Llama3 Life Interview
Extraversion 0.58 0.75 0.72 .64 .43 .45
Agreeableness 0.53 0.56 0.60 .41 .36 .35
Conscientiousness 0.59 0.66 0.63 .46 .40 .52
Neuroticism 0.63 0.59 0.75 .63 .57 .68
Openness 0.80 1.27 0.62 .43 .39 .39
Table 1: MAE and $r$ of Human Response and LLM Response Across PSI GPT-4o, PSI Llama3, and Life Interview (see Park et al., 2024, page 35 Table 3) for BFI-2 Each Personality Domain
Note: n = 357 for PSI method; n = 1,052 for Life Interview method. The bold text highlights the aspects where the PSI method performs better than the Life Interview method.
5.2 Human Personality Distribution Simulation Results
Descriptive Statistics
We compared the means and standard deviations of the different methods at three levels: personality item, facet, and domain. MAE and $r$ for both the means and standard deviations are shown in Table 2. The MAE for the means reflected the average difference between the LLM responses and the human responses at each level, whereas the MAE for the standard deviations revealed the difference in variability between the two datasets. The $r$ further illustrated the linear relationship between the two datasets, with values closer to one indicating a stronger correlation. Detailed means and standard deviations for human responses and LLM responses at the item, facet, and domain levels are provided in Tables 1213141516, and 17 in Appendix I.1.
Table 2 shows that the PSI method demonstrates better performance in simulating human samples compared to the Persona and Shape methods at the item level. Although at the facet and domain level, some results from other methods slightly outperform PSI (e.g., Shape GPT-4o in M), the process of aggregating scores from item to facet to domain levels reduces the impact of extreme values, smoothing them out at higher levels. Therefore, the PSI method’s advantage at the more granular item level becomes more important.
Additionally, PSI consistently demonstrates statistically significant positive correlations with the SD of human samples across items, facets, and domains, whereas the other methods exhibit either negative or non-significant correlations. Simulating the SD is a challenge for the other methods. Although the Shape method increases response variability and introduces more variance, it still struggles to replicate the true variance observed in human samples. The PSI method proposed in this paper addresses this issue to a certain extent.
Model Item Level Facet Level Domain Level
MAE $r$ MAE $r$ MAE $r$
M SD M SD M SD M SD M SD M SD
Persona
GPT-4o 0.48 0.44 .50 -.13 0.42 0.36 .64 -.24 0.42 0.30 .69 -.64
Llama3 0.51 0.26 .57 .02 0.33 0.15 .80 .31 0.31 0.06 .80 .76
Shape
GPT-4o 0.58 0.16 .38 -.06 0.50 0.18 .82 -.41 0.46 0.21 .84 -.77
Llama3 0.62 0.25 .45 -.26 0.52 0.38 .78 -.64 0.45 0.31 .82 -.65
PSI
GPT-4o 0.48 0.34 .68 .27 0.40 0.22 .71 .49 0.38 0.15 .67 .69
Llama3 0.66 0.36 .50 .14 0.54 0.27 .57 .19 0.49 0.21 .54 .51
Table 2: MAE and $r$ for BFI-2 Human Responses and Different Methods LLM Responses at Item, Facet, and Domain Levels
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis. The bold text indicates the best-performing data in that column.
Psychometric Performance
Psychometric performance includes multiple components; here, we primarily present model fit and structural validity (i.e., factor loadings). For other aspects, such as scale reliability and discriminant validity, please refer to Appendix I.1.
Model Fit: For both the human sample criterion and different methods of LLM responses, the TFM was fitted to each BFI-2 domain, and the FFM was fitted to all the data. TFM fit information is shown in Table 18, and FFM fit information can be seen in Table 19 in Appendix I.1.
From Table 18, it can be observed that the model fit indices of the Persona method and the PSI method are relatively similar to those of the human sample, while the Shape method performs slightly worse. Table 19 reflects a similar trend, where the Shape method shows notable discrepancies compared to the human sample in the RMSEA and SRMR model fit indices (see Appendix G for the explanation of RMSEA, SRMR, and other model fit indices).
Structural Validity: TCC was used to evaluate the similarity of factor loadings between human responses and LLM responses. TCC results for the TFM of each BFI-2 domain are shown in Table 20, while results for the FFM are shown in Table 21 in Appendix I.1.
The values of the TCC are generally high and thus supportive of factor-loading correspondence, with all methods showing strong alignment with the human sample (a TCC above .95 indicates good similarity, while a TCC of .85 to .94 suggests fair similarity; Lorenzo-Seva and Ten Berge, 2006). However, it is important to note that the TCC focuses on overall profile similarity, largely ignoring differences in absolute values (which might be too lenient). Therefore, we also need to examine the results of the specific standardized factor loadings.
Tables 22 and 23 in Appendix I.1 present the standardized factor loadings. From these tables, it is clear that the PSI method outperforms the Shape method in terms of similarity to human samples, with its factor loadings more closely matching human data. However, its results are relatively close to those of the Persona method, yet there are clearly differences in some factor loadings compared to the human sample.
It is important to note that the highest factor loadings do not necessarily indicate the best performance, as we are evaluating the similarity between the factor loadings of the human data and the LLM responses. Higher factor loadings suggest stronger correlations between the factor and the items, but this is not always ideal. It indicates that LLMs may struggle to make refined distinctions among items within a facet, meaning they may have difficulty determining which items are more or less related to the factor, unlike human respondents, who exhibit greater differentiation.
Tables 24 and 25 in Appendix I.1 present the specific inter-factor correlations. Aligning with our previous findings, the PSI method generally performs better than the other two methods when simulating human personality distributions, especially at a more fine-grained level. For instance, in the TFM results, other methods encountered anomalies, such as inter-factor correlations exceeding one (accompanied by a warning message when fitting the model), but the PSI method generally performed well.
The other results (e.g., scale reliability and discriminant validity) can be found in Appendix I.1. Overall, the results produced by the PSI method generally outperform those of the Persona method and the Shape method. However, there is still a certain gap compared to the human sample.
5.3 Personality-Related Behavioral Performance Results
Table 28 presents the correlations between the personality dimensions and OCB/CWB reported by PSI GPT-4o, with a comparison to human self-reported data. The relevant results for PSI Llama3 are provided in Table 29 in Appendix I.2, while Figures 3 and Figure 4 display the complete correlation matrices for GPT-4o and Llama3, respectively.
The results demonstrate that the correlations simulated by LLMs between OCB/CWB and various personality dimensions closely resemble the patterns seen in human self-reported data, except for the openness domain. When human reports show positive, negative, or no correlation, the LLM simulations generally exhibit consistent trends in the corresponding directions.
However, the correlations generated by LLM simulations are often higher, likely because the model primarily relies on the “typical” or “idealized” knowledge structures absorbed from its training corpus during simulation, consistently repeating and amplifying such associations. In contrast, human self-reported data contains more random noise, leading to relatively weaker correlations.

### v2

@@ 4.1 General Settings
Personality Scale
We used the Big Five Inventory–2 (BFI-2; Soto and John, 2017b), a widely used and well-validated psychological measure specifically designed to assess three distinct facets within each of the five major personality domains of the FFM. Each facet is measured by four items, resulting in a total of 60 items. Both human respondents and LLMs were explicitly instructed to rate the extent to which they agreed with each item using a 5-point Likert scale (1 = “Strongly disagree”, 2 = “Somewhat disagree”, 3 = “Neither agree nor disagree”, 4 =“Somewhat agree”, 5 = “Strongly agree”). Detailed item content and scoring criteria are presented in Table 2-4 in Appendix A.
@@
We evaluated seven widely used LLMs: Mistral-7B (Jiang et al., 2023), Gemma-2-9B and Gemma-2-27B (GemmaTeam, 2024), Llama3-8B and Llama3-70B (AI@Meta, 2024), GPT-4o-mini (OpenAI, 2024a), and GPT-4o (OpenAI, 2024b). To ensure reproducibility, all models were consistently tested at a temperature of zero. Detailed prompt descriptions are in Appendix D.
Metrics
The evaluation metrics vary across experiments and are introduced in detail within their respective sections (§4.2, §4.3, and §4.4).
@@ 4.2 Response Similarity
The main purpose of this experiment is to assess the similarity between LLM-simulated responses based on the PSI method and human self-reports.
@@
To evaluate the degree of similarity between the outputs of LLMs and human respondents, we report the Mean Absolute Error (MAE) and the Pearson correlation coefficient ($r$). These two metrics are widely used in personality modeling and psychological assessment due to their ability to capture both absolute differences and linear relationships. Specifically, lower MAE values and higher $r$ values indicate a stronger alignment between the two data sources. The formula used for calculating $r$ is provided in Appendix E.
Results
Figure 2 illustrates the $r$ between PSI method LLM-simulated data and human self-reported data. Across models, average $r$ range from 0.43 to 0.52, indicating a statistically significant and moderately strong positive association between simulated and human responses.
Figure 2: Correlation of human vs. LLM-simulated data across different LLMs and two methods. PSI achieves comparable correlations to Life Interview using only one-quarter of the time, demonstrating higher efficiency. For specific $r$ see Table 12 in Appendix I.1.
We further compared our results with prior work by Park et al. (2024), who employed a Life Interview simulation—widely regarded as one of the most information-rich methods for simulating personality data. Their work is notable for offering one of the few direct, one-to-one performance comparisons between LLMs and human participants. Compared to the Life Interview approach, the PSI method demonstrates better performance across nearly all personality domains. When evaluated using the same backbone model, GPT-4o, PSI yields a lower MAE in four domains and a higher $r$ in three. Even where PSI shows slightly higher MAE or lower $r$, the differences are minor (e.g., MAE for Openness: 0.80 vs. 0.62; $r$ for Neuroticism: 0.63 vs. 0.68; and Conscientiousness: 0.46 vs. 0.52).
Notably, the Life Interview method described by Park et al. (2024) requires approximately two hours of interview time. In contrast, the PSI approach achieves comparable outcomes with an average duration of just 34 minutes, about one quarter of the time. This significant reduction in data collection time not only lessens participant burden but also conserves computational resources needed for simulation. These results highlight the efficiency of the PSI method and its effectiveness in leveraging LLMs to simulate personality data.
@@ 4.3 Human Personality Distribution Simulation
The primary goal of this experiment is to further examine differences among prompt-based methods for simulating human personality distributions. Specifically, we compared the PSI method, the Persona method (which utilizes dialogue-based information; Zhang et al., 2018), and the Shape method (which employs adjective-based dimensional categorization; Serapio-García et al., 2023), to assess how each simulates a representative human sample.
@@
The human samples used for this experiment were collected as part of a broader project related to personality assessment through Prolific (Institutional Review Board approval was obtained). Participants were instructed to complete a set of demographic questions, the BFI-2, and a set of criterion measures. In total, 1,559 respondents provided valid responses. The average age of the participants was 42.29 years, with a standard deviation of 11.79 years, 50.80% identifying as men, 49.20% identifying as women.
Figure 3: HAI results for human vs. LLM-simulated sample across seven backbone LLMs and three methods. Each subplot shows results at a different granularity level. Higher HAIs (larger area) indicate closer alignment with the heterogeneity distribution in the human sample. PSI method performance far exceeds that of Persona and Shape.
@@
To evaluate the similarity in distribution between the human sample and the LLM-simulated sample, we leveraged multiple metrics. We computed the sample means ($\mu$) and standard deviations ($\sigma$) at the domain, facet, and item levels. For the domain and facet levels, we also calculated Cronbach’s alpha coefficients to assess internal consistency, as well as inter-scale correlations to examine the relationships among constructs. We further computed the Heterogeneity Alignment Index (HAI [^5] The HAI can be understood as a profile correlation between different observations or subgroups of $\sigma$. It has been widely applied across various fields to evaluate the similarity between samples or individuals based on multidimensional features (e.g., Humbad et al., 2013; McCrae, 2008).) by correlating the $\sigma$ profiles at each level, providing a measure of similarity in the variability structure of personality data between the human and LLM samples.
$\text{HAI}=$
$\frac{\sum_{i=1}^{n}(\sigma_{\text{H},i}\!-\!\bar{\sigma}_{\text{H}})(\sigma_{\text{M},i}\!-\!\bar{\sigma}_{\text{M}})}{\sqrt{\sum_{i=1}^{n}(\sigma_{\text{H},i}\!-\!\bar{\sigma}_{\text{H}})^{2}}\!\cdot\!\sqrt{\sum_{i=1}^{n}(\sigma_{M,i}\!-\!\bar{\sigma}_{M})^{2}}},$
@@
where $\sigma_{\text{H},i}$ and $\sigma_{\text{M},i}$ denote the $\sigma$ of the $i$-th human and LLM sample, respectively. Additionally, we conducted a three-factor confirmatory factor analysis (CFA; Jöreskog, 1969) for both the human and LLM-simulated samples, separately within each BFI-2 domain. In these models, referred to as the Three-Factor Models (TFMs), facets were treated as latent factors. Beyond modeling the facet structure within each domain, we also employed the facet-level scores as observed indicators in higher-order CFA models to evaluate the broader FFM structure. We compared model fit indices, factor loadings, and inter-factor correlations across human and LLM-simulated samples. To quantify the personality construct structural similarity between the two sample types, we used Tucker’s congruence coefficient (TCC; Tucker, 1951) and the MAE of the factor loadings. Detailed descriptions of the CFA model specifications, fit indices, and the TCC formula are provided in Appendix E. This evaluation framework is adaptable and can be applied to assess the fidelity of other types of simulated psychometric data (see Appendix H).
Descriptive Statistics Results
Figure 3 presents the HAI results comparing human sample vs. LLM-simulated samples using seven different LLMs and three simulation methods (Persona, Shape, and PSI). The three radar plots represent HAI scores at varying levels of granularity (item, facet, and domain). Each axis corresponds to a different LLM, and higher HAI values (indicated by a larger area within the plot) reflect a closer alignment with the heterogeneity patterns observed in human responses. Across all three levels of granularity, the PSI method consistently outperforms both Persona and Shape, demonstrating substantially greater alignment with human-like heterogeneity.
[…]
The above results reflect analyses of $\mu$ and $\sigma$ for both human and LLM-simulated samples across different levels of granularity. Detailed $\mu$ and $\sigma$ values are reported in Tables LABEL:tab:Mistral_dLABEL:tab:GPT-4o in Appendix I.2.
Personality Data Distribution Visualization
As shown in Figure 4, we visualized and compared personality data from the human sample and three LLM-based simulation methods across levels by projecting them into a two-dimensional space using principal component analysis (PCA). We fitted the two-dimensional PCA model using only the human sample, which served as the reference for subsequent projections of the simulated data.
Figure 4: PCA comparison of human and GPT-4o simulated data across different levels and methods. PSI distributes closest to human; Persona centers with reduced variance; Shape disperses at the extremes.
Among the three methods, PSI yielded the closest alignment with the human distribution, suggesting a stronger capacity to replicate the multivariate structure of real personality data. In contrast, the Persona method exhibited a centrally clustered distribution, indicating limited heterogeneity. The Shape method showed a more dispersed distribution, reflecting greater variability, as well as a larger divergence from the human pattern.
Model Method SOCTCC $\uparrow$ ASSTCC $\uparrow$ ENETCC $\uparrow$ SOCMAE $\downarrow$ ASSMAE $\downarrow$ ENEMAE $\downarrow$
Mistral-7B Persona $0.91$ $0.92$ $0.98$ $0.09$ $0.08$ $0.04$
Shape $0.99$ $0.94$ $0.95$ $0.03$ $0.07$ $0.07$
PSI (ours) $1.00$ $0.96$ $0.99$ $0.02$ $0.07$ $0.02$
Gemma-2-9B Persona $1.00$ $0.97$ $1.00$ $0.01$ $0.05$ $0.02$
Shape $0.97$ $0.86$ $0.97$ $0.04$ $0.10$ $0.07$
PSI (ours) $1.00$ $0.97$ $0.86$ $0.01$ $0.05$ $0.09$
Gemma-2-27B Persona $1.00$ $0.97$ $0.98$ $0.02$ $0.06$ $0.05$
Shape $1.00$ $0.97$ $0.97$ $0.02$ $0.06$ $0.08$
PSI (ours) $1.00$ $0.92$ $0.98$ $0.02$ $0.09$ $0.04$
Llama3-8B Persona 0.78-0.78 $0.55$ $0.71$ $0.41$ $0.16$ $0.14$
Shape 0.81-0.81 $0.79$ 0.80-0.80 $0.40$ $0.11$ $0.33$
PSI (ours) $1.00$ $0.98$ $0.98$ $0.02$ $0.04$ $0.04$
Llama3-70B Persona $0.99$ $0.98$ $0.98$ $0.03$ $0.05$ $0.05$
Shape $0.99$ $0.97$ $0.96$ $0.05$ $0.05$ $0.06$
PSI (ours) $1.00$ $0.95$ $0.99$ $0.03$ $0.09$ $0.03$
GPT-4o-mini Persona $1.00$ $0.99$ $0.99$ $0.01$ $0.04$ $0.05$
Shape $0.91$ $0.95$ $1.00$ $0.09$ $0.07$ $0.03$
PSI (ours) $1.00$ $0.99$ $0.98$ $0.02$ $0.03$ $0.05$
GPT-4o Persona $1.00$ $0.97$ $0.98$ $0.01$ $0.06$ $0.05$
Shape $1.00$ $0.99$ $0.97$ $0.02$ $0.05$ $0.06$
PSI (ours) $1.00$ $0.98$ $0.98$ $0.02$ $0.04$ $0.06$
Table 1: TCC and MAE for TFM of Extraversion for human vs. LLM-simulated data across different Backbone Model + Method combinations. Best TCC and MAE are in bold; Higher TCC ($\uparrow$) and lower MAE ($\downarrow$) are better.
Factor Analysis Results
Factor analysis encompasses multiple components; here, we present model fit and structural validity results (e.g., fit indices and factor loadings). Additional aspects, such as scale reliability and discriminant validity, are reported in Appendix I.2.
Model Fit: Model fit information for the TFMs and the FFM is presented in Table LABEL:tab:modelfit in Appendix I.2. For both TFMs and FFM, the model fit indices for the Persona and PSI methods are relatively comparable to those observed in the human sample, whereas the Shape method exhibits notably poorer fit. Detailed explanations of model fit indices are provided in Appendix E.
Structural Validity: TCC and MAE are used to assess the overall similarity and discrepancy in factor loadings. Table 1 illustrates the TFM results for the Extraversion domain as an example, where SOC, ASS, and ENE represent the three facet-level factors. Following Lorenzo-Seva and Ten Berge (2006), a TCC above 0.95 indicates good similarity, while values between 0.85 and 0.94 suggest moderate similarity. TCC values were generally high, indicating good alignment in factor loadings between the LLM-simulated and human samples across most methods. However, notable exceptions were observed for the Persona and Shape methods when using Llama3-8B as the backbone model, where TCC values were negative, suggesting a failure to replicate the human factor structure. In terms of MAE, the PSI method generally yielded lower errors, demonstrating better performance in capturing the true factor loading patterns. Detailed values of TCC and MAE for all TFMs and FFM are provided in Tables LABEL:tcc_tfm_e33 in Appendix I.2. Standardized factor loadings for all TFMs and the FFM are shown in Tables LABEL:tab:fc_sociabilityLABEL:tab:fc_creative_imagination and Tables LABEL:tab:ffm_extraversionLABEL:tab:ffm_openness, respectively. These patterns are consistent with the example presented in Table 1.
Tables LABEL:tab:inter_extraversion59 in Appendix I.2 report the inter-factor correlation. Consistent with our previous findings, the PSI method generally outperforms the other two methods. For example, in the TFM results, the Persona and Shape methods exhibited anomalies, such as inter-factor correlations exceeding one and accompanied by warning messages during model fitting. In contrast, the PSI method produced more stable and plausible results.
Other results like scale reliability and discriminant validity are in Appendix I.2. Overall, the PSI method consistently yields better performance compared to the Persona and Shape methods.
@@ 4.4 Personality-Related Behavioral Performance
This experiment aims to explore whether LLMs, when assigned specific personality settings, exhibit behaviors that align with theoretical expectations.
@@
We examined well-studied workplace behaviors: organizational citizenship behavior (OCB) and counterproductive work behavior (CWB), both of which are closely linked to personality (e.g., Organ and Ryan, 1995; Berry et al., 2007). Human responses were collected using the scale developed by Spector and Fox (2010), and the same items were administered to the LLMs to generate simulated data. For details on the measurement, see Appendix A; for prompt construction, refer to Appendix D.
@@
Our primary focus is on the $r$ between personality domains from different data sources and two behavioral outcomes: OCB and CWB. We expect that the $r$ between personality and OCB/CWB in the LLM-simulated data will closely mirror those observed in the human sample.
Results
Table 67 and 68 in Appendix I.3 reports the $r$ between personality domains and OCB/CWB in both human data and data simulated using the PSI method. The results demonstrate that correlations $r$ generated by PSI method largely reflect the patterns found in human data across most personality domains, with the exception of Openness. Generally, when human data exhibit positive, negative, or negligible $r$, the LLM-simulated results tend to align in the same direction. However, these simulated $r$ are often stronger in magnitude. This amplification may stem from the models’ reliance on “typical” or “idealized” knowledge structures internalized during training, which can reinforce and magnify common associations. In contrast, human responses are shaped by individual variability and random noise, leading to more diffuse patterns.
Figure 5: Ablation test results for predicting Conscientiousness. The x-axis represents each question removed on its own, and the y-axis shows the $R^{2}$ value. The red dashed line indicates the baseline $R^{2}$ when all questions are included.
4.5 Ablation Study on PSI Questions
We further investigated the individual contributions of each PSI question to personality prediction through an ablation study. The PSI was designed as a highly integrated framework, with its questions intentionally co-developed and structurally interlinked to capture diverse facets of personality-related information. As a result, isolating individual questions is inherently challenging, as doing so could undermine the integrity of the overall design.
Moreover, each question tends to tap into different aspects of personality, rendering them functionally complementary. This also implies that certain questions may be more effective than others in capturing information relevant to specific domains; some items may be more closely aligned with one domain than others. Given this design, we expected the differences in individual question effectiveness to be relatively minor overall. However, we also anticipated that the relative importance of each question would vary depending on the specific personality domain being predicted.
We conducted an ablation study by leaving one question out each time and simulated personality traits based on the remaining 31 question–answer pairs (see Table 7 in Appendix B for all questions). We then calculated the coefficient of determination ($R^{2}$) using human self-reported personality scores as the reference, in order to assess the impact of each question on model performance. If $R^{2}$ increases after a particular question is removed, this suggests that the question may have negatively affected prediction—potentially introducing redundancy, noise, or irrelevant information that weakened overall performance. Conversely, if $R^{2}$ decreases after removal, it indicates that the question made a substantial contribution to prediction and was a key factor supporting model performance.
As shown in the Figure 5, the removal of most questions leads to a slight decrease in the $R^{2}$, as one would expect, although the changes are minor and generally remain within a ±0.01 range. Notably, the relative importance of individual questions varies across different prediction domains (see Figure 7-10 in the Appendix I.4). This empirical finding supports our earlier hypothesis: different types of questions capture distinct facets of personality, offering complementary rather than redundant information. Furthermore, the relative importance of questions varies across target domains, indicating that different questions make distinct contributions to domain-specific predictions. Together, these findings underscore the meaningful contributions of each PSI question.